Operations Report – January 27th

SOL: 6
Name of person filing report: Scott Dorrington
Non-nominal systems: None
Notes on non-nominal systems: None
Curiosity rover used:
Hours: 264.9
Beginning charge: 100%
Ending charge: 68%
Currently charging:
Perseverance rover used:
Hours: 282.8
Beginning charge: 99%
Ending charge: 62%
Currently charging: No
Spirit rover used:
Hours: 243.4
Beginning charge: 100%
Ending charge: 43%
Currently charging: No
Opportunity rover used:
Hours: 157.8
Beginning charge: 100%
Ending charge: 51%
Currently charging: No
General notes on rovers: None
Summary of Hab operations:
WATER USE: 26.1 gallons
Water (static tank): 331.1 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): on
Toilet tank emptied: Yes
Summary of internet: Nominal during comms window
Summary of suits and radios:
Suit #8 still stood down. Located some replacement nuts for charging port, but not correct size. Will continue repair tomorrow.
Summary of GreenHab operations: put notes here
WATER USE: 5 gallons
Heater: On automatic
Supplemental light: on at 6pm – to be fixed to turn on at 10pm
Harvest: 6g basil, 5g parsley, 1g sag, 17g arugula
Summary of ScienceDome operations: Alex entered to observe crew of EVA #10 upon their return.
Dual split: Nominal
Summary of RAM operations: (Tools used, work done) put notes here
Summary of any observatory issues: See astronomy report
Summary of health and safety issues: None.
Questions, concerns and requests to Mission Support:
Toilet seat has been replaced. Thank you.
May have follow up questions on the greenhab heater/thermostat in afternoon pending observations.

Copyright © The Mars Society. All rights reserved. | Main Site